One number is six less than the other. The sum of the squares of these numbers is 20. Find the numbers.
Answer by checkley71(8403)

X=Y-6
X^2+Y^2=20
(Y-6)^2+Y^2=20
Y^2-12Y+36+Y^2=20
2Y^2-12Y+36-20=0
2Y^2-12Y+16=0
2(Y^2-6Y+8)=0
2(Y-4)(Y-2)=0
Y-4=0
Y=4 ANSWER. & X=4-6 OR X=-2
Y-2=0
Y=2 ANSWER. & X=2-6 OR X=-4